# Social Media Isn’t Just Instagram: A Youth-Envisioned Platform for Meaningful Social Connections

### v1

@@ 4.5.2. Reduced Celebrity and Commercial Influence
Other participants expressed a need for control over how non-personal content is organized and placed on the platform. P04 suggested being able to “figure out how I want to arrange strangers’ spaces in my… platform.” They also wanted the ability to “prioritize things and feel more in control of how much information is fed to them,” imagining a system where apps and notifications are ranked by personal relevance. P18 saw ads as inevitable but wished they could appear in a separate “billboard” that users could skip.
4.5.3. Spatial Representation of Relationships
Youth envisioned SMH as a platform that reflects the dynamics of real-world relationships, using physical distances within the SMH landscape to intuitively represent the strength of social ties. Participants appreciated this design as it prioritized interactions with closer friends, creating a sense of community and authenticity. P02 highlighted, “The fact that only my close friends have access to it… is the defining factor,” enabling users to focus on “better conversations with other people [that] already know who I am,” without the pressure to “advertise who I am” as they might feel on MSM.
Some participants envisioned SMH landscape to resemble a college campus or a neighborhood, where physical closeness naturally fosters interaction. P03 imagined, “The people around you… would be people from the same house as you,” cultivating a shared sense of “togetherness” and “community building” through common spaces and interactions. Similarly, P01 proposed a tiered layout, where “the top 7 people you’re closest to would be the 7 houses in your cul-de-sac area,” while acquaintances would be positioned farther away, requiring a “little hike” to reach them.
This design approach—aligning physical landscapes with emotional closeness—was seen as a way to streamline interactions and focus on meaningful relationships. P07 noted how, in short bursts of social media use, they preferred seeing content from their closest friends over “some girl I talked to once three years ago.” Unlike MSM feeds, which participants felt prioritized engagement over relationship strength, SMH’s spatial arrangement allowed users to engage naturally with those closest to them. P09 further elaborated on how this physical representation could enhance navigation: “have your friends’ rooms, or like your kind of close friends list, and maybe you’re kind of living in a pod with them, so you can move between the rooms a lot quicker.”
4.5.4. Intentional and Meaningful Use of Time
Participants expressed a strong preference for digital experiences that feel meaningful and intentional, contrasting these with the negative feelings often associated with aimless social media use. P10 described how social media can feel “meaningless” when they “didn’t learn something,” while P18 reflected on the regret they feel after “mindlessly scrolling.” Similarly, P14 shared that scrolling through Instagram Reels while “literally just laying in [their] bed” feels like “wasting [their] time.”
In contrast, participants found interactions involving direct communication and engagement more fulfilling. P16 shared that “anytime not scrolling” and instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this sentiment, noting that social media use feels more meaningful when they are “actually talking to someone.” They explained that such interactions create a sense of “connectivity,” particularly when keeping in touch with people “not in close proximity.”
The distinction between passive and active engagement was another recurring theme. P10 explained that while doomscrolling does not make them feel “fulfilled,” playing games like Minecraft is different because it involves “progress” they can “save and come back to.” Similarly, P14 shared that the lack of connectivity on social media drives their “doomscrolling,” but they believe platforms that encourage interaction could reduce screen time by making time spent online more meaningful. P06 also note that they noted that they “don’t feel bad… most of the time [they] use Roblox,” as it involves engaging activities such as “playing games with friends” and “customizing spaces.”
Given such differences in how time spent is perceived, participants expected time spent on SMH to feel more meaningful than that on MSM. P04 noted that while MSM allows for quicker interpersonal interactions, it often leads to “doomscrolling.” In contrast, SMH would, for example, encourage “thoughtfully engaging with someone’s memories,” leaving them feeling “much more fulfilled.” P05, similarly, explained that they would not feel bad about spending more time on a platform like SMH because it involves “more effort and more time… a more intentional way of interacting.” P07 affirmed this, stating they would be willing to spend more time on a platform like SMH given interactions where they would find interactions “more enjoyable.”
4.6. Reimagining Social Media’s Potential

### v2

@@ 4.3.3. Prioritizing Close Relationships and Reducing Extraneous Content
4.3.4. Developing Socially Grounded and Themed Conversational Spaces
Participants described how themed, socially grounded conversational spaces designed for protected, meaningful dialogue could make SMH feel more personal and emotionally connected than MSM.
“Anyone you’re talking to in general, just emphasizing conversation between individuals more, rather than just posting on a thread. Like how I talked last time about Reddit, you know, there’s a thread, and then you make a comment on it, but there are a hundred other comments, so no one’s going to respond to you. I want to focus more on having actual conversations with people, making real connections, and more individualized chatting. I find it’s way easier to get to know people a lot better.” (P22)
Participants described how the design and feel of conversational spaces could make social interactions on SMH feel safer, more personal, and more inviting than on existing platforms. Many valued the idea of walking into themed, topic-specific rooms where conversation felt protected and purposeful, rather than broadcasting to a broad, undefined audience. P22, for example, imagined “chat rooms you can walk into,” each centered on themes like “high school stuff,” decorated with artifacts from shared experiences to help spark meaningful conversations. They described these as “protected space[s]…where you can talk…without feeling unsafe or judged,” contrasting them with the more exposed, contextless feel of mainstream social media where “you don’t know everything around it.”
Participants also imagined personalizing these conversational spaces to make them feel more emotionally connected. P07 described filling chatrooms with “photos of you together, inside jokes, or fun memories,” making even casual conversations feel more intimate. P09 expanded on this idea, imagining a “Room of Requirement”-like system (roo, [n. d.]) where users could create custom rooms on demand to match the “purpose” of their conversation. Participants also valued interest-based spaces, similar to Discord (discord page, [n. d.]), where a “common point” (P13) or “good talking point” (P07) could help break the ice and give conversations direction.
4.3.5. Implementing Comprehensive Boundary Controls to Create Personal Spaces that Feel Safe
To ensure SMH feels like a safe, personal environment, participants stressed the need for more comprehensive boundary controls beyond privacy tools found on existing platforms that they perceived as limited and inconsistent. P22, for example, explained that even with contact syncing turned off on Instagram, the platform still suggested their Finsta (Fake + Instagram; secondary account (Huang and Vitak, 2022)) account to people they were trying to avoid, based on mutual follow relationships. They stressed the need for stronger, more thorough blocking controls, including the ability to block “alternate accounts” and fully remove any trace of a blocked person’s presence, explaining, “I don’t wanna be able to tell that the person I blocked is even around me.” P14 similarly pointed out that even private accounts on Instagram still expose personal content, like Story Highlights, to “a random person”, revealing perceived gaps in platform-level privacy protections (Bojkov, 2019).
Participants expressed a desire for stronger control over their own visibility, imagining features that would let them opt out of being seen or discovered entirely. P03 suggested a feature like an “invisibility cloak” for moments when users “don’t want anyone to see you.” P07 proposed a “super private” profile mode where users “can’t even be searched… like you’re not even on the app” unless they choose to make themselves visible.
Preventing unsolicited contact from strangers was another recurring priority. P22 proposed a model inspired by Discord, where only people you’ve already interacted with can contact you, and others “can’t even be discovered unless you have something in common.” This idea resonated with P13, who argued that new connections should require mutual consent, suggesting that people “should only be able to add you if you meet in person and exchange a QR code.” They shared a personal experience of being able to find a stranger online based on just their first name, reflecting that “it’s too easy to find people on social media.”
4.4. Nuanced Self-Presentation
Participants described MSM as “a popularity contest” (P22) that felt “really fake” (P14) because “people only really show the good moments” (P14). This carefully curated image culture creates “a lot of seriousness” (P23), which is reflected in how “actors or celebrities… [even] have somebody else run their Instagram account” (P23). P13 shared an anecdote about helping a friend prepare an Instagram post, including selecting the “post order,” highlighting “how seriously we take social media and our image online.” In contrast, participants imagined SMH as a space where self-presentation could feel more personal, authentic, and flexible, reflecting the full range of who they are, without the pressures of curation or performance.
4.4.1. Expressing Identity Through Customizable Avatars, Spaces, and Artifacts
Participants consistently described customizable avatars and personal spaces as meaningful tools for self-expression, offering more creativity, personality, and authenticity than the static profiles or limited settings found on MSM. Unlike MSM, which typically offers only “two modes” (P10), i.e., “light and dark” (P10), participants imagined a much richer and more immersive experience. For example, P21 envisioned a fully customizable world where users could build their own “mini world” within a larger virtual space, complete with “houses, teleport points, and rooms like a kitchen, a bathroom, a bedroom, or even a school.” These spaces could be shared with friends as a way of “showing a little part of yourself.” Similarly, P03 described wanting to create a “dark academia” Hogwarts-inspired room, filled with “so many bookshelves and… Victorian architecture,” to reflect their personal interests.
Customizable “3D avatar[s]” (P11) and profiles were also seen as key ways of showing individuality. P22 reflected on the early days of Myspace (mys, [n. d.]), describing how they enjoyed coding their profile layouts and changing “the font and really make it reflect me.” Even small details, like choosing a username, were described as ways to “convey to the world and your friends your own identity… which is a positive feeling usually.” P01 noted that such forms of expression are often “mainly for you,” not just for public display. P17 added that even if users don’t consciously intend to reveal something personal, their choices still do: “your brain conjured that up, right? So I think that’s part of you.”
Beyond visual customization, participants imagined SMH supporting more ambient, indirect forms of expression, offering subtle cues about what someone is doing, feeling, or thinking without needing to make a formal post. For instance, P01 described wanting their avatar to reflect real-time states, like the outfit they are wearing or the song they are listening to, so that friends could intuitively notice things like, “‘Oh. that’s what she’s listening to right now’ or ‘That’s what she’s thinking about.”’ Similarly, P10 imagined that friends could pick up on changes in their interests simply by observing how they update their virtual space over time, providing “an insight into… what I’m thinking.”
Several participants highlighted the contrast between the flat, compressed feel of 2D platforms and the immersive qualities of 3D environments. P10 noted that in 3D, others can see “as much stuff as you can see in 3D,” making it feel richer and more complete. P16 added that VR creates a sense of presence that feels more like meeting “real people… compared to the profile pictures,” because you can pick up on their “mannerisms” and notice details like “the way they hold themselves… the way they decorate their house… how they choose to dress their in-game character.”
Participants also described how these kinds of choices feel more authentic than generic posts because they emerge naturally from personal preferences and interests. As P01 put it, such “creative expression” feels “a lot more personal” because it reflects “that’s not just like what life has given you… a reflection of your outlook on life.” P13 explained that these subtle forms of sharing make social media feel more like “an extension of everyday interaction,” helping people “gain the basic foremost information” about someone without needing to start a direct conversation. P21 suggested this would make it “easier for people to approach you and say, ‘Hey, I like that, too.”’
Drawing parallels to the physical world, P13 pointed out how personal spaces like bedrooms carry a sense of intimacy and authenticity that translates well to SMH. They explained:
“A bedroom is a respected personal space… If you label it as a bedroom, it just feels more intimate. It’s where you spend most of your time, and it shows your personality the most. It’s not like you can easily change the stuff you have in your room.… If someone were to come over, you can’t just rearrange it in 30 minutes or 30 seconds the way you can on Instagram. If you wanted to add someone there, all you’d have to do is remove a couple of things. It’s easier than physically changing a space.” (P13)
Finally, participants appreciated these ambient and low-pressure forms of expression for their lower social stakes. P16, who described themselves as “fairly private” and not particularly “public” about their preferences, explained that “decorat[ing] a house” or customizing an avatar allows others to “get to know more about my personality… through my style” without feeling like they are performing for attention. They described this kind of expression as “more justified” and less “embarrassing” than posting content just to be seen by others.
4.4.2. Enabling Multifaceted Self-Presentation Across Spaces and Relationships
Participants envisioned SMH supporting flexible, multifaceted self-presentation, making it easier to share different sides of themselves across various spaces and relationships, and reducing the pressure to maintain a single, polished identity. Rather than being limited to a single, polished identity, as is often the case on MSM, participants wanted ways to represent the full range of their interests, identities, and relationships. P01 described this as having a house with rooms “dedicated to each of your interests,” explaining, “I’m into multiple things. You could have a room for Harry Potter, a room for neuroscience, and a room for another fandom or interest.” This ability to represent different facets of oneself was seen as a way to reduce misjudgment and give others a more complete picture of who they are. P02 built on this idea, imagining that users with multiple “Muggle personalities” could organize them into separate spaces, saying, “Each personality has its own vault.”
4.4.3. Sharing Everyday Activities and Lived Experiences to Deepend Understanding and Connection
Participants valued the ability to share casual, everyday moments and rich, embodied memories, filling the gaps between the highlights usually shared on MSM into a more natural, meaningful extension of daily life and personal connection. P04 described wanting to share small, ordinary moments, like “what I’m doing at the moment. For example, if I’m at the shops, maybe I’d want to share that.” P21 extended this idea to items purchased, such as a “new pair of New Balance shoes,” imagining virtual spaces that automatically displayed “recently added or recently stored items.”
Participants also imagined SMH supporting the sharing of more personal, embodied memories as a way to deepen understanding and reduce misinterpretation. P04 described wanting to share lived experiences like “jumping into a pool or taking the first bite of an ice cream sundae,” drawing inspiration from the Pensieve in Harry Potter (pen, [n. d.]) as a way to create stronger emotional connections. Unlike MSM, where communication often feels like “just words,” (P03) participants saw memory-sharing as a way to communicate context, emotion, and intention more fully. P02 explained that this kind of sharing could help others “understand why I said what I said” and prevent “misconstru[al]s.” (P02)
Some participants imagined that these shared memories could include sensory or emotional details, not just visuals. P01, for example, expressed interest in being able to “transmit smells through social media,” adding another layer to digital communication. Others compared it to live streaming personal experiences. P07 described it as “screen sharing, but with your eyes,” and P08 highlighted its usefulness in capturing fleeting moments, like crossing a state line, without the need to pause and record. P03 imagined being able to instantly share moments with loved ones “in the moment,” rather than having to explain them after the fact, when the emotion or relevance might already have passed.
4.4.4. Flexible, Contextual Privacy as a Foundation for Authentic Self-Presentation
Participants emphasized that flexible, contextual privacy is essential for supporting authentic self-presentation, allowing people to manage who sees what, when, and how, without feeling overexposed or misrepresented. They described current platforms as offering rigid and limited controls that fail to account for the complexity of real social life. P16 explained that on Instagram, for example, privacy options are reduced to features like “close friends,” adding, “it’s kind of a pain to organize.” P11 similarly noted that as people share “more like their genuine real life and their experiences,” privacy needs to become “a little bit more complicated,” requiring finer-grained control over visibility.
Participants also reflected on the challenge of managing different audiences and social contexts on MSM and imagined SMH supporting more seamless transitions between identities and spaces. P02 described a system that would “just telepathically know what’s in Muggle area and what’s in Hogwarts area,” allowing users to “walk through a portal” and appear as “a new person” without needing separate accounts. Building on this, participants imagined contextual identity management, where users could shift how they present themselves based on the audience or situation. P03 proposed a model where everyone maintains a “base” identity but customizes additional layers. P22 similarly described having different avatars or profiles for different groups, like on Discord, where “any group you’re in… you could change your avatar for that.”
Participants also stressed that contextual privacy should extend beyond people-based access to include temporal and relational factors. P14 proposed limiting access to different stages of one’s past identity, saying, “I’m cool with you seeing the 6-year-old version of me on this, but I don’t want you to see my 12-year-old version.” Others, like P21, suggested gradual openness, where people could selectively reveal more of themselves over time. As they described, “if you’re meeting a complete stranger, you wouldn’t want to give them like exactly how you look… But then, after a couple of conversations, you sort of get the gist of how they are, and whether you want them to see you… so then you can turn it on. Or turn it off.”
Participants also highlighted the importance of individual control over these transitions and boundaries. P14 explained that privacy settings should not be one-size-fits-all, suggesting that “each of them should have their own setting, because I feel like everybody has different levels of what they want to keep to themselves.” Together, these suggestions reflect flexible, contextual privacy as not just a technical feature, but as a core requirement for supporting the kinds of personal, meaningful, and low-pressure interactions youth hoped SMH would make possible.
4.5. Intentional Engagement
Participants described their growing frustration with how current social media platforms prioritize passive consumption over meaningful connection. In response, they envisioned SMH as a space that promotes intentional, socially rewarding interactions. This section explores four key themes that reflect their desire to refocus social media on authentic connection, purposeful engagement, and meaningful use of time.
4.5.1. Encouraging Deliberate Navigation and Content Discovery
Participants imagined SMH as a space that encourages deliberate, effortful discovery, replacing the passive, algorithm-driven feeds of MSM with experiences that require users to actively choose what and who they engage with. P22 described their frustration with Instagram, explaining that “half of the timeline is accounts I don’t even follow that were just randomly recommended to me,” turning it into “less keeping up with friends and more just getting random stuff shoved in my face.” Instead, participants hoped SMH would encourage more intentional, purposeful engagement, both when browsing content and when navigating the virtual world.
P04 described how slowing down to explore someone’s space or story could make interactions feel more personal and meaningful. Unlike Instagram, where you “see all the posts at once” and can quickly scroll past them, SMH would require users to “take more time to hone in on a memory,” offering the satisfaction of consciously consuming content from someone you’re close to, rather than just “taking a quick second to look at it… and then being done.”
Participants also imagined features that would prompt users to actively choose what they wanted to engage with before entering a space. P09 compared this to using an elevator, where you decide on your destination in advance: “Could [SMH] be sort of like that, where you think about what you want to see before entering that space, and then it shows what you were thinking about?” Others imagined brain-controlled interfaces that would surface content based on what users genuinely want in the moment, not just what an algorithm predicts will keep them engaged.
Physical movement through SMH’s world was also seen as key to this intentionality. Participants valued having to walk, explore, and navigate to find people, spaces, or activities, turning even small actions into social moments of presence and care. P03 explained that this kind of embodied movement makes SMH feel “more connected than regular social media,” (P03) because it requires active participation rather than passive scrolling.
This desire for intentional interaction extended to how users move through the platform itself. Participants imagined physical navigation, such as walking or traveling through virtual space, as a way to make engagement feel more meaningful. P02 explained that they would prefer slower, more thoughtful interactions over instant teleportation: “I wouldn’t really want anything to be super fast decisions. I want everything to be very thought through.” While features like “spells” (P06) or the “Marauder’s Map” (P21) were imagined as convenient options, participants valued the ability to walk and explore, turning simple actions like visiting a friend or browsing a shop into shared social moments. P03 explained that this kind of embodied interaction feels “more connected than regular social media” (P03) because it requires active presence rather than passive scrolling.
P05 also imagined SMH as a place where users would intentionally seek out content or people, rather than having information pushed to them. They described SMH as a world where you might “walk through a town to get there,” explaining, “It’s not challenging to navigate, but you have to intentionally go there to see it. It isn’t something you can do mindlessly.” P14 agreed, noting that visiting someone’s space or content in SMH would feel much more personal than encountering “a random 2D Insta post that they posted 3 months ago.”
Participants believed that this kind of effortful navigation would naturally discourage casual lurking or influencer-driven engagement, shifting the platform’s focus toward genuine connection. As P05 put it, “Influencers probably wouldn’t be as keen on something that is more time-consuming… People who are genuinely just trying to connect with others would probably benefit… because it is more personal.”
4.5.2. Making Social Media About (Meaningful) Connection, Not (Passive) Consumption
Beyond how users navigate content, participants also emphasized the importance of shifting the platform’s core purpose from endless consumption toward meaningful social connection. Participants expressed frustration with how current platforms promote mindless, passive consumption, and described their hopes for SMH to instead foster more meaningful, intentional connections through active social engagement. They described frustration with how easily time slips away on platforms like Instagram and TikTok without leaving them feeling unfulfilled. P10 noted that social media often feels “meaningless” when they “didn’t learn something,” and P18 reflected on the regret of “mindlessly scrolling.” P14 similarly described spending time on Instagram Reels while “literally just laying in [their] bed” as “wasting [their] time.”
A major contributor to this feeling was the infinite scrolling design of MSM. P11 explained that “there’s always more content on social media,” which fuels the fear of missing out and the urge to keep scrolling. P14 critiqued TikTok for exposing users to “clickbaity” and often low-quality content that “doesn’t stimulate [users] at all” and “messes with [their] attention span.”
In contrast, participants described active, social engagement, even if casual, as far more rewarding. P16 shared that “anytime not scrolling” but instead “talking to people” feels worthwhile, “even if it’s just light conversation.” P14 echoed this, noting that meaningful use comes when they are “actually talking to someone,” especially to stay connected with people “not in close proximity.” They reflected that while users might spend more time on a platform like SMH, that time would feel worthwhile because “at the end of the day, you’re connecting with a person that you can’t otherwise connect with… and that makes it more valuable in and of itself” compared to passively watching “50-second TikTok videos of somebody making slime for like an hour.”
Participants also contrasted passive consumption with active participation in interactive environments. P10 explained that while doomscrolling feels unproductive, activities like playing Minecraft feel different because they offer “progress” that can be “saved and come back to.” P06 echoed this sentiment, noting that they “don’t feel bad… most of the time [they] use Roblox,” because it involves “playing games with friends” and “customizing spaces.”
Reflecting on SMH’s potential to create this kind of active, engaging experience, P09 acknowledged that “time spent might not be mitigated,” but emphasized that “the way you’re spending it” would make the difference. They explained that SMH would “feel a lot more interactive instead of just taking in content, and more and more and more content,” suggesting that meaningful engagement, rather than passive consumption, was what ultimately mattered for them.
4.6. Playfulness and Whimsy
Participants described SMH as a platform that feels youthful, playful, and deeply immersive, offering not just entertainment but opportunities for creative self-expression, social bonding, and emotional well-being. By embracing whimsy and delight, SMH was seen as making social media feel more personal, less pressured, and more meaningful.
4.6.1. Embracing Creative Joy and Playful Self-Expression
Participants imagined SMH as a platform that transforms social media into a space for creative joy and playful self-expression, offering users the freedom to experiment with identity, explore personal interests, and connect with others through creativity. Several participants described how SMH offered far more than static profiles or simple feeds. P06 reflected that “everyone likes sharing parts of themselves,” appreciating how the platform allows users to shape not just what they say, but how they appear and interact with others. P11 imagined a “dream social media world” where users could fully redesign their profiles, “maybe their avatar in the middle, maybe their posts on either side, maybe a bio below”, allowing them to “express their creativity and, like, yourself” in new ways not possible on conventional platforms.
For some, this act of personal customization felt like a meaningful process of identity exploration. P01 described the deeper satisfaction in “just that process of creating it, and finding out who you are through that process.” P21 echoed this sentiment, describing SMH as feeling “more like a game” because of the “creative freedom” it provides, making the experience feel “more fictional than real” in a way that encourages experimentation and self-expression. P22 added that SMH offered something they hadn’t found anywhere else, explaining, “I can really be creative here… I can make something that I haven’t made anywhere else, you know. That’s where the fun aspect comes in, really, for me.”
This spirit of creativity extends beyond the individual to the wider environment and community. Participants described the excitement of exploring what others had created, likening it to visiting realms in Minecraft where you “get to see how like everyone puts in their own effort into it… Whoa! They built that, that’s so cool” (P06). SMH felt alive and engaging whether users were socializing or spending time alone: “it’s not just like, Oh, the group chats are only fun when there’s other people here. No, it’s fun when we’re by ourselves, too, we get to do our own things” (P06). The interactive environment itself, where “you’re not just sitting there chatting, you’re also walking around the space… you’re interacting with whatever else is in that area” (P06), offered more to do than conventional platforms. Whether through building, exploring, or engaging with shared content, SMH transformed social interaction into something playful, creative, and deeply engaging.
4.6.2. Designing Engaging and Playful Privacy Mechanisms
Participants imagined privacy not as a tedious and confusing task but as part of the playful experience, envisioning interactive, game-like mechanisms that turn boundary-setting into something fun, discoverable, and socially meaningful. P11 envisioned systems where users could control access to spaces through gestures like a “secret knock” or hidden objects that revealed “secret passages” accessible only to trusted friends. Others suggested knowledge-based gates, such as “secret handshakes” (P21) shared among friends. Drawing inspiration from the Marauder’s Map, P02 imagined privacy controls activated by “spells”, unlocking spaces only when correctly summoned.
These features were described not just as functional but as adding delight and novelty to social interaction. P05 compared them to “little Easter eggs,” remarking that “there’s that little kid part of everyone” that makes such features enjoyable. They contrasted these with conventional privacy settings on mainstream social media, which they described as “a wall of settings” that often feel confusing or uninviting.
4.6.3. Augmenting Everyday Interactions with Fantastical Interfaces and Environments
Participants imagined SMH using magical and fantastical technologies to transform everyday digital interactions, making them feel more immersive, accessible, and emotionally uplifting, even for those with limited real-world opportunities. P01 questioned the need for phones at all, asking, “If you’re already carrying around like a wand everywhere with you, why would you also want to carry a phone?” They envisioned wands projecting “holographic screens, like Jarvis in Iron Man,” or activating “panels” and “big screens” through spells. Others suggested alternative magical devices. P15 preferred VR glasses, describing them as “more natural,” “easier,” and “more fun.” Participants also imagined objects like a “makeup mirror” that could serve as a portal, a “watch” that immerses users into SMH, or a “mirror glass” for video calls or summoning friends with spells like accio.
These magical affordances were not only imagined as novel interfaces but also as ways to make extraordinary experiences more accessible. P09 highlighted that SMH could give people “experiencing stress” or lacking access to “really, really beautiful landscape and environments” the chance to walk around and “admir[e]” calming, scenic spaces. P16 imagined that users could take part in activities they might never otherwise experience, saying, “Maybe you’re busy… you can’t afford to travel to Nepal to climb Mount Everest… but you have a day, so you can just log in for the entire day and just climb a mountain.”
4.7. Expanding Horizons